x cubed - 9x = 0
and
x squared + x - 2 = 0

Answer by checkley71(8403)

X~3-9X=0 OR X(X~2-9)=0 OR X(X-3)(X+3)=0 X=3 & X=-3
X~2+X-2=0 OR (X-1)(X+2)=0 X=1 & X=-2